# Synthetic Turbulence via an Instanton Gas Approximation

### v1

@@ I Introduction
Figure 1: Three independent snapshots of random velocity field realizations $u(x)$, $x\in[0,2\pi]$, for each of the six random field ensembles analyzed in Sec. VII (columns), for three different noise variances $\sigma^{2}\in\{1.17\cdot 10^{-1},1.85\cdot 10^{2},2.51\cdot 10^{4}\}$ (rows), corresponding to Reynolds numbers $\text{Re}\in\{1.03,32.29,192.21\}$ for DNS of the Burgers Eq. (1). Technical details are explained in Sec. II to VI, and statistical results are analyzed in Sec. VII. On a high level, we can see how the Burgers DNS fields (column 1 from the left) display stronger and stronger negative gradients or quasi-shocks, as well as smooth ramps, as the Reynolds number increases. This is reproduced to varying degrees by the instanton gas ensembles (columns 2 to 4), but not by the structureless synthetic fields (columns 5 and 6).
@@
As for potential limitations of the instanton-gas method, we first point out that recent advances in machine learning provide an alternative, and practically successful, route to turbulence generation, e.g. via generative adversarial networks [53, 54] or diffusion models [55, 56, 57, 58]. Nevertheless, we aim for a more physically principled and interpretable synthetic turbulence simulation here, which retains its merits despite those advances. Furthermore, computing the necessary instanton configurations for our method numerically through optimization can remain a technical challenge to implement. However, recently, progress has been made to develop efficient and manageable approaches, see e.g. Refs. [59, 60, 61, 62]. Lastly, as we discuss in more detail below, there are a few to a certain extent arbitrary – but physically motivated – parameters in the method implemented in this paper, such as the choice of observable, and the number of instantons for the canonical ensemble, that may require some fine-tuning. Still, compared to sampling structured random fields using other parametric families of fields with similar methods as discussed in Sec. IV, the instanton-based approach could be regarded as more canonical, as it removes a lot of the arbitrariness and reduces the number of free parameters.
@@
The outline of the paper is as follows. First, we briefly recall the necessary elements of the instanton formalism to calculate turbulence statistics in Sec. II. Numerical details and results of DNS of the Burgers Eq. (1), as well as the single-instanton formalism in itself, are presented in Sec. III. Subsequently, we derive the instanton gas approximation used for synthesizing turbulent fields in Sec. IV. We will then discuss our Monte Carlo approach to sample from different instanton gas ensembles in Sec. V, where the main ingredients are the one-point probability density function (PDF) $\rho(a)=\mathbb{E}\left[\delta(\partial_{x}u(x_{0},t_{0})-a)\right]$ of the gradients $\partial_{x}u$ in the turbulent field, a Gram determinant $g$, and possible interactions $\Delta S$ between instantons which enter the steps of a Markov chain Monte Carlo (MCMC) algorithm. Other random field ensembles without coherent structures for comparison are briefly introduced in Sec. VI. We will then present numerical results for the different random field ensembles on Eulerian statistics — PDFs of gradients, energy spectra, and structure functions — as well as particle statistics, which exhibit similarities to the characteristics of cosmic ray propagation in turbulent magnetic fields, in Sec. VII. We conclude with an outlook on the necessary steps to extend this method to MHD turbulence, ultimately enabling its application to cosmic ray propagation, in Sec. VIII. Appendix A contains additional details about the selection of hyperparameters for the log-normal cascade model.
@@ II The instanton formalism
$C(a)&=\lvert\lambda_{a}\rvert\left[2S^{\text{I}}(a)\operatorname{det}\left(\operatorname{Id}-B_{a}\right)\right]^{-1/2}\\ &=\lvert\lambda_{a}\rvert\left[2S^{\text{I}}(a)\prod_{k=1}^{\infty}\left(1-\mu_{a}^{(k)}\right)\right]^{-1/2}\,.\end{aligned}$ (5)
@@ III Numerical details for instanton computations and DNS
Figure 2: Reynolds number Re as a function of the noise variance $\sigma^{2}$ for DNS of the Burgers Eq. (1), as detailed in Sec. III. The scaling line $\text{Re}\propto\sigma^{2/3}$ corresponds to the theoretically expected results at large $\sigma$ through rescaling the system [90, 78]. For all further results shown in this paper, we will report $\text{Re}=L_{\text{box}}u_{\mathrm{rms}}/\nu$ (orange squares) to label the different simulations.
Figure 3: Numerically computed instanton action $S^{\text{I}}$ (top) and 1-loop PDF prefactor $C$ (bottom) from Eq. (6) as a function of the gradient value $a=\partial_{x}u(0,t_{0})$. For numerical details, see Sec. III. Each open dot in the figures corresponds to one instanton configuration. Both axes are logarithmically scaled, except for the shaded region around $a=0$ where the horizontal axis is linear. Scaling lines are shown for comparison; the ones for the action at large positive and negative gradients values are in accordance with the literature [45, 47, 91, 90].
As a prerequisite for all subsequent instanton-based synthetic turbulence sampling strategies, we numerically calculate and store instanton configurations $u^{\text{I}}_{a,0}(x,t)$ for about $500$ values of $a\in[-30000,-0.1]$ and about $70$ values of $a\in[0.1,223]$, both logarithmically spaced, with an optimization code that employs a pseudo-spectral discretization and a second-order explicit Runge–Kutta time stepper with integrating factor for the diffusion terms, cf. Ref. [59]. We use a spatial resolution of $n_{x}=512$ for smaller gradient strengths $a$ and $n_{x}=2048$ for the largest ones, as well as a temporal resolution of $n_{t}=10^{4}$ in the time interval $[-T,t_{0}]=[-1,0]$ for all calculations. The size of the time interval has been chosen such that DNS statistics are approximately stationary at the final time $t_{0}=0$. The “Mexican hat” forcing correlation in Fourier space reads
@@
Figure 4: Numerically computed instanton configurations $u^{\text{I}}_{a,0}$ at final time $t=0$ (top) and in space and time (bottom) for three different observable values $\partial_{x}u(0,0)=a\in\{-255,-12,11\}$ (columns). For negative gradient values, we see a shock-like structure that steepens and becomes more localized in time as the absolute value of the prescribed gradient at the final time at the origin increases. For positive gradient values, we see a smoother ramp-like structure. For numerical details of the optimization procedure and parameters used, see Sec. II and III, as well as Ref. [91] for a detailed discussion of the time evolution of instantons of Eq. (1).
@@
Figure 5: One-point gradient PDF $\rho$ at different Reynolds numbers as measured from DNS of the Burgers Eq. (1) (dots), compared to the instanton prediction with 1-loop correction $\rho^{(1)}$ according to Eq. (6) (solid lines). For a better comparison, the horizontal axis has been rescaled to the respective root-mean-square gradient strength of each DNS run, and both the DNS data and instanton prediction have also been shifted vertically by the same factor per Reynolds number. As expected, we see good agreement for small to moderate Re, while at least the scaling of the instanton prediction is still correct at large Re. For comparison, we also show the left tail of $\rho^{(0)}(a)=\text{const}\cdot\exp\left\{-S^{\text{I}}(a)/\sigma^{2}\right\}$ (dashed lines), i.e. the theoretical instanton prediction for the one-point gradient PDF using just the instanton, with an unknown (and hence assumed to be constant) prefactor. Evidently, including the Gaussian fluctuations improves the agreement considerably, and removes the unknown normalization constant.
@@
We show the Reynolds numbers Re that corresponds to the DNS runs at different $\sigma^{2}$ in Fig. 2. Here, we computed the root-mean-square velocity $u_{\text{rms}}=\mathbb{E}\left[u(\cdot,t_{0})^{2}\right]^{1/2}$, energy $E=\tfrac{1}{2}u_{\text{rms}}^{2}$, the mean dissipation density $\epsilon=2\nu\mathbb{E}\left[\partial_{x}u(\cdot,t_{0})^{2}\right]$, and the integral length $L_{\text{int}}=E^{3/2}/\epsilon$ from the DNS data at final time for each noise variance $\sigma^{2}$. We see that the DNS runs cover about three decades of Re. Next, we show the results of the instanton and Gaussian fluctuation computations for the necessary quantities $S^{\text{I}}(a)$ and $C(a)$ to approximate the one-point gradient PDF in Fig. 3. We remark that the power-law scaling of the prefactor $C(a)$ over a wide range of gradient strengths could be the subject of a future theoretical study. Plots of the computed shock or ramp-like instanton configurations $u^{\text{I}}_{a,0}(x,t)$ for a few different values of $a$ are shown in Fig. 4.
@@
Finally, by combining the instanton action $S^{\text{I}}(a)$ and 1-loop prefactor $C(a)$ according to Eq. (11), we obtain the prediction for the one-point gradient PDF $\rho$ at different $\sigma^{2}$ shown in Fig. 5 (solid lines), which compares them to the gradient PDFs as measured in the DNS runs. We see excellent quantitative agreement for small to moderate noise variances $\sigma^{2}$. As expected, for larger $\sigma^{2}$ the agreement becomes worse, even though the scaling of the DNS PDFs is still captured correctly by the “instanton-plus-Gaussian-fluctuations” approximation. Importantly, as noted in Ref. [78] already, we see that the 1-loop prefactor correction $C(a)$ in $\rho^{(1)}(a)$ is in fact quite important for this agreement: The dashed lines $\propto\rho^{(0)}(a)$ in Fig. 5 highlight that the instanton action, without accounting for fluctuations, indeed only describes the far left tail of the PDF [90, 84].
@@ IV.1 Canonical ensemble of instantons
where []delimited-[][$\dots$][ … ] stands for repeating the preceding expression. Here, the additional Jacobian term from the noise-to-velocity transformation [93, 94] in the action at finite $\sigma$, proportional to the functional divergence of the deterministic terms in the equation of motion $u\partial_{x}u-\partial_{xx}u$, can be neglected [95], e.g. because the instantons for the Burgers Eq. (1) are all antisymmetric functions in space with respect to $x_{0}$. Now, restricting the fields to lie on the submanifold described by Eq. (8) leads to
@@
×$\times\exp\left\{-\frac{1}{\sigma^{2}}S\left[u\left(\left\{a_{j},x_{0j}\right\}_{j=1}^{n_{\text{I}}}\right)\right]\right\}$ (11)
@@
in this case. Adding and subtracting the single-instanton actions in the exponent of Eq. (11), we can alternatively write
@@
×$\times\exp\left\{-\frac{1}{\sigma^{2}}\Delta S\left(\left\{a_{j},x_{0j}\right\}_{j=1}^{n_{\text{I}}}\right)\right\}$ (14)
@@
IV.2 Grand canonical ensemble of instantons
It is intuitive to expect that if – instead of prescribing a fixed number of instantons $n_{\text{I}}$ in Eq. (8) – the number of instantons $n_{\text{I}}$ is allowed to vary, with a self-consistently determined distribution, then a better approximation to the true random field ensemble could be obtained. This corresponds to using a grand canonical ensemble of instantons in the language of statistical mechanics. If fluctuations in the number of instantons around $\mathbb{E}\left[n_{\text{I}}\right]$ are not negligible, we may expect it to possibly behave differently from a canonical ensemble. The relative weight and hence probability of an $n_{\text{I}}$-instanton configuration is determined by the $n_{\text{I}}$-instanton partition function. Typically, interacting particles in a gas within the grand canonical ensemble are simulated by adding random particle addition and removal moves to the Metropolis sampler of Sec. V [98, 52]. For the present work, we leave this is a future direction to explore, and only consider canonical ensembles, with an instanton number $n_{\text{I}}=3$ throughout, chosen heuristically in line with DNS snapshots.
However, we remark that in the interacting canonical instanton ensemble of the previous subsection, since an instanton $u^{\text{I}}_{a_{j},x_{0j}}$ effectively “disappears” from Eq. (8) as $a_{j}\to 0$, it is possible that due to the interactions, the ensemble also assigns a non-negligible probability to configurations with effectively fewer than $n_{\text{I}}$ instantons. One should hence think of the interacting canonical ensemble considered here as prescribing an upper bound $n_{\text{I}}$ on the number of instantons, but still allowing for fewer instantons to be present in principle. In contrast to this, the non-interacting canonical ensemble (16) strictly fixes the number of instantons $n_{\text{I}}$.

### v2

@@ I Introduction
Figure 1: Three independent snapshots of random velocity field realizations $u(x)$, $x\in[0,2\pi]$, for each of the six random field ensembles analyzed in Sec. VII (columns), for three different noise variances $\sigma^{2}\in\{1.17\cdot 10^{-1},1.85\cdot 10^{2},2.51\cdot 10^{4}\}$ (rows), corresponding to Reynolds numbers $\text{Re}\in\{1.03,32.29,192.21\}$ for DNS of the Burgers Eq. (1). Technical details are explained in Sec. II to VI, and statistical results are analyzed in Sec. VII. The Burgers DNS fields (column 1) display stronger and stronger negative gradients or quasi-shocks, as well as smooth ramps, as the Reynolds number increases. This is reproduced to varying degrees by the instanton gas ensembles (columns 2 to 4), but not by the structureless synthetic fields (columns 5 and 6).
@@
As for potential limitations of the instanton-gas method, we first point out that recent advances in machine learning provide an alternative, and practically successful, route to turbulence generation, e.g. via generative adversarial networks [53, 54] or diffusion models [55, 56, 57, 58]. Nevertheless, we aim for a more physically principled and interpretable synthetic turbulence simulation here, which retains its merits despite those advances. Furthermore, computing the necessary instanton configurations for our method numerically through optimization can remain a technical challenge to implement. However, recently, progress has been made to develop efficient and manageable approaches, see e.g. Refs. [59, 60, 61, 62]. Lastly, as we discuss in more detail below, there are a few to a certain extent arbitrary – but physically motivated – parameters in the method implemented in this paper, such as the choice of observable, and the number of instantons for the canonical ensemble, that may require some fine-tuning. Still, compared to sampling structured random fields using other parametric families of fields with similar methods as discussed in Sec. IV, the instanton-based approach could be regarded as more canonical, as it removes a lot of the arbitrariness and reduces the number of free parameters. However, we would like to point out that an instanton gas approximation of higher-dimensional and complex systems still raises many open questions. The question of suitable observables must be physically motivated. In the case of the Burgers equation discussed next, this is natural, but in the case of the three-dimensional Navier–Stokes equations, for example, the question must then determine whether vorticity or strain is chosen. In the example of MHD turbulence shown in the outlook, the current density is chosen as the observable, since an ensemble of current sheets plays a central role in the diffusion of cosmic rays.
@@
The outline of the paper is as follows. First, we briefly recall the necessary elements of the instanton formalism to calculate turbulence statistics in Sec. II. Numerical details and results of DNS of the Burgers Eq. (1), as well as the single-instanton formalism in itself, are presented in Sec. III. Subsequently, we derive the instanton gas approximation used for synthesizing turbulent fields in Sec. IV. We will then discuss our Monte Carlo approach to sample from different instanton gas ensembles in Sec. V, where the main ingredients are the one-point probability density function (PDF) $\rho(a)=\mathbb{E}\left[\delta(\partial_{x}u(x_{0},t_{0})-a)\right]$ of the gradients $\partial_{x}u$ in the turbulent field, a Gram determinant $g$, and possible interactions $\Delta S$ between instantons which enter the steps of a Markov chain Monte Carlo (MCMC) algorithm. Other random field ensembles without coherent structures for comparison are briefly introduced in Sec. VI. We will then present numerical results for the different random field ensembles on Eulerian statistics — PDFs of gradients, energy spectra, and structure functions — as well as particle statistics, which exhibit similarities to the characteristics of cosmic ray propagation in turbulent magnetic fields, in Sec. VII. Among the different possible instanton gas ensembles that are introduced theoretically in Sec. IV, the ones that we sample numerically for the results of this paper in Sec. IV are: a canonical ensemble of $n_{\text{I}}=3$ non-interacting instantons without fluctuations around them (ensemble 2), a canonical ensemble of $n_{\text{I}}=3$ interacting instantons, without approximations to their interaction $\Delta S$, and without fluctuations around them (ensemble 3), and a canonical ensemble of $n_{\text{I}}=3$ non-interacting instantons with Gaussian fluctuations around the individual instantons (ensemble 4). We do not study simplified interaction models for $\Delta S$ or grand-canonical ensembles of instantons numerically in this paper, leaving this as future work. We conclude with a summary and discussion in Sec. VIII, and provide an outlook on the necessary steps to extend this method to MHD turbulence, ultimately enabling its application to cosmic ray propagation, in Sec. IX. Appendix A contains additional details about the selection of hyperparameters for the log-normal cascade model.
@@ II The instanton formalism
$\begin{aligned} C(a)&=\lvert\lambda_{a}\rvert\left[2S^{\text{I}}(a)\operatorname{det}\left(\operatorname{Id}-B_{a}\right)\right]^{-1/2}\\ &=\lvert\lambda_{a}\rvert\left[2S^{\text{I}}(a)\prod_{k=1}^{\infty}\left(1-\mu_{a}^{(k)}\right)\right]^{-1/2}\,.\end{aligned}$ (5)
@@ III Numerical details for instanton computations and DNS
Figure 2: Reynolds number Re as a function of the noise variance $\sigma^{2}$ for DNS of the Burgers Eq. (1). The dahsed line $\text{Re}\propto\sigma^{2/3}$ corresponds to the theoretically expected scaling at large $\sigma$ [90, 78]. For all further results, we report $\text{Re}=L_{\text{box}}u_{\mathrm{rms}}/\nu$ (orange squares) to label the simulations.
Figure 3: Numerically computed instanton action $S^{\text{I}}$ (top) and 1-loop PDF prefactor $C$ (bottom) from Eq. (6) as a function of the gradient $a=\partial_{x}u(0,t_{0})$. Each open dot corresponds to one instanton configuration. Both axes are logarithmically scaled, except for the shaded region around $a=0$ where the horizontal axis is linear. Scaling lines are shown for comparison; the ones for the action at large positive and negative gradients values are in accordance with the literature [45, 47, 91, 90].
As a prerequisite for all subsequent instanton-based synthetic turbulence sampling strategies, we numerically calculate and store instanton configurations $u^{\text{I}}_{a,0}(x,t)$ for about $500$ values of $a\in[-30000,-0.1]$ and about $70$ values of $a\in[0.1,223]$, both logarithmically spaced, with an optimization code that employs a pseudo-spectral discretization and a second-order explicit Runge–Kutta time stepper with integrating factor for the diffusion terms, cf. Ref. [59]. We use a spatial resolution of $n_{x}=512$ for smaller gradient strengths $a$ and $n_{x}=2048$ for the largest ones, as well as a temporal resolution of $n_{t}=10^{4}$ in the time interval $T,t_{0}]=[-1,0]$ for all calculations. The size of the time interval has been chosen such that DNS statistics are approximately stationary at the final time $t_{0}=0$. The “Mexican hat” forcing correlation in Fourier space reads
@@
Figure 4: Numerically computed instanton configurations $u^{\text{I}}_{a,0}$ at final time $t=0$ (top) and in space and time (bottom) for three different observable values $\partial_{x}u(0,0)=a\in\{-255,-12,11\}$ (columns). For negative gradient values, we see a shock-like structure that steepens and becomes more localized in time as the absolute value of the prescribed gradient at the final time at the origin increases. For positive gradient values, we see a smoother ramp-like structure.
@@
Figure 5: One-point gradient PDF $\rho$ at different Reynolds numbers from DNS of the Burgers Eq. (1) (dots), compared to the instanton prediction with 1-loop correction $\rho^{(1)}$ in Eq. (6) (solid lines). The horizontal axis has been rescaled to the respective root-mean-square gradient strength of each DNS run, and both the DNS data and instanton prediction have been shifted vertically by the same factor per Reynolds number for better visulatization. We also show the left tail of $\rho^{(0)}(a)=\text{const}\cdot\exp\left\{-S^{\text{I}}(a)/\sigma^{2}\right\}$ (dashed lines), i.e. the theoretical instanton prediction using only the instanton, with an unknown (and hence assumed to be constant) prefactor.
Figure 6: Local scaling exponent $\vartheta$, defined in Eq. (8), of the one-point gradient PDFs from Fig. 5 for negative gradients $a<0$. The colors are the same as in Fig. 5, with blue meaning small Re, and red large Re. The dashed black line is the theoretical prediction using only the instanton, while the solid lines show the (Re-dependent) predictions using the instanton and Gaussian fluctuations. For the latter, some numerical artifacts from inaccuracies in the instanton optimization and prefactor computation, amplified through the use of finite differences for $\mathrm{d}{\cal S}/\mathrm{d}a$, are visible.
@@
We show the Reynolds numbers Re that corresponds to the DNS runs at different $\sigma^{2}$ in Fig. 2. Here, we computed the root-mean-square velocity $u_{\text{rms}}=\mathbb{E}\left[u(\cdot,t_{0})^{2}\right]^{1/2}$, energy $E=\tfrac{1}{2}u_{\text{rms}}^{2}$, the mean dissipation density $\epsilon=2\nu\mathbb{E}\left[\partial_{x}u(\cdot,t_{0})^{2}\right]$, and the integral length $L_{\text{int}}=E^{3/2}/\epsilon$ from the DNS data at final time for each noise variance $\sigma^{2}$. We see that the DNS runs cover about three decades of Re. Next, we show the results of the instanton and Gaussian fluctuation computations for the necessary quantities $S^{\text{I}}(a)$ and $C(a)$ to approximate the one-point gradient PDF in Fig. 3. We remark that the power-law scaling of the prefactor $C(a)$ over a wide range of gradient strengths could be the subject of a future theoretical study. Plots of the computed shock or ramp-like instanton configurations $u^{\text{I}}_{a,0}(x,t)$ for a few different values of $a$ are shown in Fig. 4. We refer to Ref. [91] for a detailed discussion of the time evolution of instantons for the Burgers Eq. (1).
@@
By combining the instanton action $S^{\text{I}}(a)$ and 1-loop prefactor $C(a)$ according to Eq. (12), we obtain the prediction for the one-point gradient PDF $\rho$ at different $\sigma^{2}$ shown in Fig. 5 (solid lines), which compares them to the gradient PDFs as measured in the DNS runs. We see excellent quantitative agreement for small to moderate noise variances $\sigma^{2}$. As expected, for larger $\sigma^{2}$ the agreement becomes worse, even though the scaling of the DNS PDFs still appears to be captured correctly by the “instanton-plus-Gaussian-fluctuations” approximation. Importantly, as noted in Ref. [78] already, we see that the 1-loop prefactor correction $C(a)$ in $\rho^{(1)}(a)$ is crucial for this agreement: The dashed lines $\propto\rho^{(0)}(a)$ in Fig. 5 highlight that the instanton action, without accounting for fluctuations, indeed only describes the far left tail of the PDF [90, 84].
For a more quantitative analysis of the one-point gradient PDF scaling in Fig. 5, we consider the local scaling exponent [91, 90]
$\vartheta(a):=\frac{\mathrm{d}\log{\cal S}(a)}{\mathrm{d}\log\lvert a\rvert}=\frac{a}{{\cal S}}\frac{\mathrm{d}{\cal S}}{\mathrm{d}a}\,.$ (8)
Here, we define $\cal S$}, the exponential part of the PDF $\rho$, as follows: For either the DNS results or any of the PDF approximations through instantons at a given $\sigma^{2}$, we write $\rho(a)=c\exp\{-{\cal S}(a)/\sigma^{2}\}$. The constant $c=c(\sigma^{2})$ is fixed through the condition ${\cal S}(0)\overset{!}{=}0$  [^1] Otherwise one could add any additive constant to $\cal S$} and compensate for it in $c$, rendering $\vartheta$ not uniquely defined, such that ${\cal S}(a)=-\sigma^{2}\log(\rho(a)/\rho(0))$. If the local scaling exponent $\vartheta(a)$ has a plateau somewhere, then the PDF follows a stretched exponential distribution $\propto\exp\{-\beta\lvert a\rvert^{\vartheta}\}$ with exponent $\vartheta$ in that region. This is of particular interest, since such distributions are known to occur for extreme events in other turbulent systems as well [93]. If we approximate the PDF using just the instanton without fluctuations via $\rho^{(0)}$, we have ${\cal S}(a)=S^{\text{I}}(a)$, and by the sensitivity property of Lagrange multipliers, the predicted exponent is $\vartheta(a)=\lambda_{a}a/S^{\text{I}}(a)$ for any $\sigma^{2}$. For the 1-loop approximation $\rho^{(1)}$ in Eq. (6) including Gaussian fluctuations, we have ${\cal S}(a)=S^{\text{I}}(a)-\sigma^{2}\log(C(a)/C(0))$, and hence there is a $\sigma^{2}$-dependent tradeoff between the action and prefactor. For both $\rho^{(1)}$ and the DNS PDFs, we use first-order finite differences to approximate the derivative in Eq. (8). Since the DNS data is noisy, in particular in the tails, we subsequently apply a smoothing filter to the resulting $\vartheta(a)$ for a cleaner visualization. The result of this procedure is shown in Fig. 6 for the more interesting case of negative gradients. The figure confirms that (i) for large $\lvert a\rvert$, the prefactor always becomes subdominant at any $\sigma^{2}$, and the scaling exponent hence asymptotically approaches the instanton prediction $\vartheta(a)=a\lambda_{a}/S^{\text{I}}(a)$ as $\lvert a\rvert\to\infty$, (ii) the instanton prediction without fluctuations interpolates non-monotonically between Gaussian scaling $\vartheta=2$ for small $\lvert a\rvert\ll 1$, which follows from linearizing Eq. (3), and $\vartheta=3/2$ asymptotically as $a\to$-$\infty$ [47], and (iii) for finite $\lvert a\rvert$ and high $\sigma^{2}$, the local scaling exponent in the DNS data can be very far from $a\lambda_{a}/S^{\text{I}}(a)$, but still turns out to be well-described by the scaling exponent of $\rho^{(1)}$. Point (iii) confirms our prior observation that the scaling of the DNS PDFs is captured correctly using the instanton and Gaussian fluctuations around it, even at the highest Reynolds number considered here. The mismatch in the absolute normalization of the PDFs, which leads to the constant offset observed at high Re in Fig. 5, could hence be viewed as an artifact of misrepresenting the core of the PDF at high Re.
@@ IV.1 Canonical ensemble of instantons
where $\dots$] stands for repeating the preceding expression. Here, the additional Jacobian term from the noise-to-velocity transformation [95, 96] in the action at finite $\sigma$, proportional to the functional divergence of the deterministic terms in the equation of motion $u\partial_{x}u-\partial_{xx}u$, can be neglected [97], e.g. because the instantons for the Burgers Eq. (1) are all antisymmetric functions in space with respect to $x_{0}$. Now, restricting the fields to lie on the submanifold described by Eq. (9) leads to
@@
$\hskip-14.22636pt\times\exp\left\{-\frac{1}{\sigma^{2}}S\left[u\left(\left\{a_{j},x_{0j}\right\}_{j=1}^{n_{\text{I}}}\right)\right]\right\}$ (12)
@@
in this case. Note that the Gramian $g$, despite appearing as a pre-exponential factor in the PDF (12), is purely geometric in nature, and does not correspond to any random fluctuations around the instanton fields.
Adding and subtracting the single-instanton actions in the exponent of Eq. (12), we can alternatively write
@@
$\times\exp\left\{-\frac{1}{\sigma^{2}}\Delta S\left(\left\{a_{j},x_{0j}\right\}_{j=1}^{n_{\text{I}}}\right)\right\}$ (15)
@@
IV.2 Grand canonical ensemble of instantons and the choice of the number of instantons $n_{\text{I}}$
It is intuitive to expect that if – instead of prescribing a fixed number of instantons $n_{\text{I}}$ in Eq. (9) – the number of instantons $n_{\text{I}}$ is allowed to vary, with a self-consistently determined distribution, then a better approximation to the true random field ensemble could be obtained. This corresponds to using a grand canonical ensemble of instantons in the language of statistical mechanics. If fluctuations in the number of instantons around $\mathbb{E}\left[n_{\text{I}}\right]$ are not negligible, we may expect it to possibly behave differently from a canonical ensemble. The relative weight and hence probability of an $n_{\text{I}}$-instanton configuration is determined by the $n_{\text{I}}$-instanton partition function. Typically, interacting particles in a gas within the grand canonical ensemble are simulated by adding random particle addition and removal moves to the Metropolis sampler of Sec. V [100, 52].
For the present work, we leave this is a future direction to explore, and only consider canonical ensembles, with an instanton number $n_{\text{I}}=3$ throughout, chosen heuristically in line with DNS snapshots (see the discussion below). However, we remark that in the interacting canonical instanton ensemble of the previous subsection, since an instanton $u^{\text{I}}_{a_{j},x_{0j}}$ effectively “disappears” from Eq. (9) as $a_{j}\to 0$, it is possible that due to exclusion-type interactions, the ensemble assigns a non-negligible probability to configurations with effectively fewer than $n_{\text{I}}$ instantons. One should hence think of the interacting canonical ensemble considered here as prescribing an upper bound $n_{\text{I}}$ on the number of instantons, but potentially resulting in fewer instantons being present with high probability. In contrast to this, the non-interacting canonical ensemble (17) strictly fixes the number of instantons to $n_{\text{I}}$. In the latter ensemble, a smaller effective number of instantons in a given random field sample can e.g. result from independently drawing two instanton positions that are close to the same location, and with comparable absolute value of their gradient strengths of either the same or opposite sign.
We can estimate the typical spatial extent of a single coherent structure for the Burgers Eq. (1) and the corresponding instanton (not only the shock width itself, but the entire structure) as the correlation length $L_{\chi}$ of the large-scale forcing covariance $\chi$ (in the present setting: $L_{\chi}=1$, $L_{\text{box}}=2\pi$). Hence, the number of instantons $n_{\text{I}}$ should be chosen as $L_{\text{box}}/L_{\chi}$ up to a dimensionless constant of $O(1)$. Choosing a canonical ensemble of instantons with $n_{\text{I}}\ll L_{\text{box}}/L_{\chi}$ will lead to a dilute, weakly interacting gas, whereas $n_{\text{I}}\gg L_{\text{box}}/L_{\chi}$ leads to strong interaction effects. In general, a theoretical determination of $n_{\text{I}}$ for the canonical ensembles would involve calculating or approximating $\mathbb{E}[n_{\text{I}}]$ within a grand-canonical one [49]. In contrast to this, for this paper, in order to determine $n_{\text{I}}$ for numerical simulations of canonical ensembles, we counted peaks in $\partial_{x}u$ snapshots from DNS at different Re, which indicated that no more than $n_{\text{I}}=4$ instantons should be used. We then repeated the simulations and (Eulerian) statistical analysis shown in Sec. VII for $n_{\text{I}}=1,\dots,4$, and present results for the representative value $n_{\text{I}}=3$ for all Re in this paper, which gave the closest overall visual and statistical agreement with DNS.